OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (3, –1) and is perpendicular to the line y = x + 1?

OpenStudy (anonymous):

y = x + 1 has slope 1, so a perpendicular to it will have slope -1 From the general equation for a line you have: y = -x + b, where b is the intercept, you substitute 3 and -1 in it and you get: -1 = -3 + b, so b = 2 Your line is y = -x + 2